A SPRING HAVING A CONSTANT OF 44N/M IS INITIALLY COMPRESSED 4.0CM IT IS THEN COMPRESSED ANOTHER 5.0CM DETERMINE THE WORK DONE TO COMPRESS THE SPRING FROM 4.0CM TO 9.0CM
1
Expert's answer
2019-12-12T06:18:11-0500
"W=\frac{kx_2^2}{2}-\frac{kx_1^2}{2}=\frac{k}{2}(x_2^2-x_1^2)""=\frac{44 \frac{N}{m}}{2}((0.09 m)^2-(0.04 m)^2)""=0.143 J"